OpenStudy (anonymous):

Volume of a gas varies directly as the temperature. If V=3 when T=150, Find V when T is 200.

OpenStudy (anonymous):

The answer is 4, right? @ganeshie8

ganeshie8 (ganeshie8):

\(\large \dfrac{3}{150} = \dfrac{V}{200}\) \(\large \implies \dfrac{3}{150} \times 200 = V\)

ganeshie8 (ganeshie8):

simplifying, gives u \(V = 4\) you're right ! good job !!

OpenStudy (anonymous):

Yaaaaaaaay!! lol Thank you ! :)

ganeshie8 (ganeshie8):

np :)